# LLaMA-Berry: Pairwise Optimization for O1-like Olympiad-Level Mathematical Reasoning

### v1

@@ 3.3 Ablation Study
Table 4: Performance Breakdown of LLaMA-3.1-8B-Instruct with SR-MCTS
GSM8K MATH500 AIME2024
Zero-shot CoT 76.6 46.2 6.7
One-turn Self-Refine 84.4 55.4 10.0
Rollouts or Samples 8 16 8 16 8 16
Major Voting with Random Repeated Sampling 88.3 88.9 46.8 48.2 10.0 10.0
Self-Consistency 88.5 88.6 43.8 43.0 10.0 10.0
Tree of Thoughts 83.2 82.5 40.2 35.0 6.7 10.0
Multi-turn Self-Refine with Self-Evaluation 78.9 78.2 40.2 40.4 6.7 6.7
CoT based MCTS with Process Self-Reward 68.5 70.7 33.2 33.8 6.7 3.3
Self-Refine based MCTS with Outcome Self-Reward 82.0@rm8 81.4@rm16 37.2@rm8 36.4@rm16 6.7@rm8 6.7@rm16
SR-MCTS (Self-Refine applied to MCTS with PPRM) 95.0@rm8 96.1@rm16 69.0@rm8 76.6@rm16 16.7@rm8 26.7@rm16
In our research, we design a series of ablation experiments based on the concept of a contrast matrix, systematically evaluating the key components of our methods across three benchmarks of increasing difficulty: GSM8K, MATH500, and AIME2024. We incorporate three baseline methods as control groups: Zero-shot CoT, One-turn Self-Refine, and Major Voting with Random Repeated Sampling. By comparing the performance of our experimental groups with these baselines, we assess the necessity of adopting more sophisticated mechanisms.
Self-Refine based Search vs. CoT based Search: In the GSM8K and MATH500 benchmarks, the Self-Refine method demonstrates significant advantages over Zero-shot CoT. Specifically, in the GSM8K task, Self-Refine achieves accuracy rates of 95.0% and 96.1% for 8 and 16 rollouts, respectively, substantially surpassing the 76.6% achieved by CoT. In the MATH500 task, the difference is even more pronounced, with Self-Refine achieving accuracies of 69.0% and 76.6% for 8 and 16 rollouts, compared to CoT’s 46.2%.
Self-Reward vs. PPRM: Comparing the self-reward mechanism with PPRM, the latter exhibit particularly strong performance in the GSM8K benchmark. The SR-MCTS using PPRM achieves a high accuracy of 96.1% with 16 rollouts, significantly better than the process and outcome self-reward approaches, which only reach 81.4% and 70.7%. PPRM, by considering both the process and performance comprehensively, more effectively motivates the model to explore optimal solutions, thereby significantly enhancing performance.
These findings underscore the efficacy of combining the Self-Refine method with PPRM when addressing complex problems. Lastly, the contrast between self-reward mechanisms and PPRM emphasizes the importance of designing reward mechanisms that consider both process and outcome comprehensively. PPRM provides a more holistic incentive to the model, thus fostering more effective problem-solving strategies.
@@ 3.4 Scaling Study
Figure 4: Scaling of Test-time Rollouts
Scaling of Test-time Rollouts: To explore the potential and trends of the scaling with rollouts in test-time, we depict the pass rate with rollouts in three benchmarks with different difficulty levels. Analyzing the performance alongside the Figure 4, the increment in the number of rollouts consistently enhances model performance across various benchmarks, and the extent of these improvements differs depending on the benchmark’s complexity and the base model’s reasoning capability. The LLaMA-3-8b-Instruct model displays a steady progression in performance with each increase in rollouts but faces more pronounced challenges in the complex MATH500 benchmark, suggesting a limitation in handling intricate mathematical reasoning without substantial computational support. In contrast, the LLaMA-3.1-8b-Instruct model begins with a higher baseline performance across all benchmarks. This indicates that its self-refine capabilities better utilize the additional computational resources provided by more extensive rollouts. This is particularly evident in the MATH500 benchmark, where the performance disparity between the two models significantly expands, highlighting the critical role of foundational model capabilities in managing sophisticated tasks.
Figure 5: Average Token Consumption Comparison Across Datasets
Figure 6: Average Token Consumption for LLaMA-3.1-8B-Instruct Across Olympiad-level Datasets
Scaling of Test-time Token Overheads: Comparing LLaMA-3-8B-Instruct and LLaMA-3.1-8B-Instruct across GSM8K, GSMHARD, and MATH500, LLaMA-3.1-8B-Instruct consistently consumes more tokens across all categories—Solutions, Critiques, and Overall. This 5 result reflects LLaMA-3.1-8B-Instruct’s tendency to generate more detailed and comprehensive outputs with increased overhead. While this likely improves solution quality, it also introduces greater resource demands and variability in token usage, highlighting a trade-off between accuracy and computational resources. LLaMA-3.1-8B-Instruct is thus better suited for tasks prioritizing precision over speed. As shown in 6, Token overheads during inference also scale with task difficulty across different olympiad-level benchmarks with LLaMA-3.1-8B-Instruct. AIME2024 exhibits the highest token consumption with significant variability, reflecting the complexity of its solution paths. In contrast, GPQA Diamond shows lower overall overhead, while AMC2023 falls in between, with moderate token consumption and less variability than AIME2024 but still notable.
4 Related Work
4.1 PRM&ORM
The mathematical reasoning capabilities of LLMs exhibit continuous enhancement. OpenAI o1 (OpenAI, 2024) even achieves ability comparable to graduate students on challenging GPQA benchmarks. Recent studies mainly focus on collecting high quality data or domain-specific knowledge (Liao et al., 2024; Huang et al., 2024; Toshniwal et al., 2024; Lu et al., 2024; Yu et al., 2023b; Zhao et al., 2024) or training external reward models (Kang et al., 2024; Wang et al., 2023; Havrilla et al., 2024; Lightman et al., 2023; Ma et al., 2023). There are primarily two types of reward models. The first one is Outcome-supervised Reward Models (ORMs), which are trained from the final CoT answer of the model, while ORMs assign a reward to a complete answer. This kind of feedback is not precise enough, and it is challenging to align the reward. The second one is process-supervised reward models (PRMs), which receive feedback at each step in CoT. PRMs are considered to be more effective (Lightman et al., 2023). However, PRMs face significant challenges such as the need for extensive manually annotated data (Luo et al., 2024a; Havrilla et al., 2024), and the inefficiency in guiding LLMs to generate answers (Feng et al., 2023; Ding et al., 2023). To address these issues, we propose the Pairwise Preference Reward Model (PPRM), which uses the following instructions and the learning ability of LLMs in context. Our approach effectively improves the efficiency and accuracy of answer search by directly comparing the quality of two answers and globally ranking all generated answers. The whole process saves the cost of collecting much human-annotated data.
4.2 Tree Search Reasoning
@@
Sampling diverse reasoning paths (Brown et al., 2024) can significantly enhance the probability of finding the correct answers. Self-Consistency (Wang et al., 2022) sample a complete path each time while tree search methods like Tree-of-Thought (ToT) (Yao et al., 2023) and MCTS (Chen et al., 2024a; b; Luo et al., 2024b; Feng et al., 2023; Xie et al., 2024; Xu, 2023; Liu et al., 2023; Tian et al., 2024; Ding et al., 2023) extend numerous steps to optimize step answers and ultimately obtain the optimal solution. Additionally, Self-refine (Madaan et al., 2023b) method has become a recent focus. Self-verification (Gero et al., 2023; Weng et al., 2022) and rStar (Qi et al., 2024) utilize the inherent capabilities of the model to iteratively explore and refine answers. However, the performance of self-refine is typically constrained by the inherent capabilities of the model, especially for small language models (SLMs) with significantly weaker self-refine abilities (Madaan et al., 2023a). Our approach explores the answers by combining MCTS with self-refine. Ultimately, we construct a global win-loss matrix of answers following a directed graph process.
@@ 5 Conclusion
The research aims to address the challenges in complex mathematical reasoning, especially at the Olympiad level, where generating accurate and efficient reasoning paths for problem-solving. By combining iterative optimization methods Self-Refine applied to Monte Carlo Tree Search (SR-MCTS), the LLaMA-Berry framework improves the efficiency of solution generation. The Pairwise Preference Reward Model (PPRM) further enhances performance by modeling preferences between solutions rather than just scoring outcomes. Experimental results demonstrate that LLaMA-Berry outperforms baseline approaches on benchmarks like GSM8K and MATH, and demonstrates a competitive level with GPT-4 Turbo on difficult benchmarks without additional training.
In practice, the LLaMA-Berry framework still faces several limitations. The potential high computational cost of Monte Carlo Tree Search (MCTS) and Self-Refine methods may limit their practicality in resource-constrained environments. Additionally, the framework’s assessment has been still narrow, focusing primarily on mathematical reasoning benchmarks, and its effectiveness on larger parameter or close-source models remains unexplored, warranting future experiments to assess its scalability and broader applicability.
@@ Appendix A Hyper-paramerer settings
All hyper-parameter settings for Section 3 are listed in Table 5.
Table 5: Hyper-parameter settings
@@ Appendix B Details of Grading and Metrics
Grading: We follow the correctness evaluation method proposed in PRM800K (Lightman et al., 2023) to score the answers generated by the algorithm. For the mathematical solutions proposed by the algorithm and their corresponding ground truth, we inform the model of the expected response format in a prompt. The answer’s formula or value is extracted by matching the response format with predefined rules. If the model fails to follow the expected format in the prompt and the rule-based extraction fails, the answer is directly judged as inconsistent.
@@
    Following the criterion proposed in PRM800K. (Lightman et al., 2023), we use the Sympy library to simplify the difference between the expression corresponding to the answer label, denoted as $a$, and the expression corresponding to the ground truth label, denoted as $b$. If the simplification yields $a-b\iff 0$, the criterion is satisfied.
Metrics: To provide a comprehensive and robust evaluation metric, we adopt major@k and pass@k based on reward-ranked sorting (denoted as rm@k) as the evaluation metrics.
  • pass@k represents the fraction of tasks that pass evaluation for at least one of the top $k$ samples generated. This is calculated using the unbiased estimator:
[…]
    where $n\geq k$ is the number of samples generated per task, $c\leq n$ is the number of correct samples that pass the unit tests, and $\binom{n}{k}$ represents the binomial coefficient. The reward-ranked metric, rm@k, follows the same definition but sorts the samples by their reward values before evaluation.
  • major@k is defined as the fraction of tasks where a majority of the top $k$ samples generated return the same correct solution. This metric focuses on consistency across multiple generated answers. Majority weight is calculated on answer’s extracted labels.
For results without a subscript, the score represents the greedy evaluation using the default prompt of the base model. Result with a notation indicate the use of the corresponding prompt engineering technique, and those marked with self-consistency use the self-consistency aggregation method.
@@ C.1 Overview
The design goal of the PPRM is to integrate the properties of both PRM and ORM, providing a more nuanced ordinal response between any two solution answers. We attempt to utilize reinforcement learning methods for training, leveraging the model’s instruction-following capability to predict the relative merits of pairs of problem-solving answers. This will further enable the use of the EBC method to evaluate the global quantile scores of mathematical solutions.
@@ C.2 Data Collection
Our data synthesis is derived from two datasets: PRM800K and OpenMathInstruct-1. The PRM800K dataset, collected from the MATH dataset, comprises a substantial number of step-divided problem-solving answers, with manual quality annotations for each step. We primarily utilize this dataset to generate pairs of answers for comparative analysis based on step-wise process quality. The OpenMathInstruct-1 dataset incorporates data from the GSM8K and MATH datasets, which have been manually annotated for outcome correctness. We use this dataset to synthesize pairs for comparative analysis based on outcome quality.
In processing PRM800K, for a given problem, we first sample steps of varying quality annotations from the step-wise dataset to construct a complete reasoning path. For pairs with the same final step annotation, paths composed of higher-quality steps are considered superior to those with lower-quality steps. In cases where the final step annotations are identical, the reasoning path with the superior final step annotation is regarded as better.
@@
Notably, we filtered out any data samples from PRM800K and OpenMathInstruct-1 that may overlap with the GSM8K and MATH test sets. Ultimately, we formed a dataset of 7,780,951 entries for training the PPRM model.
@@ C.3 Direct Preference Pair Construction
Figure 7: Dataset Construction of PPRM
For all pairs, we frame the inquiry as ”For Question $Q$, is solution $a_{1}$ better than solution $a_{2}$?” If solution $a_{1}$ is deemed superior to solution $a_{2}$, we label it as ’Yes’; otherwise, it is labeled as ’No’.
@@ C.4 RLHF Training
Figure 8: RLHF Training of PPRM
We subsequently apply the DPO method to train the Gemma2-2B-it model using RLHF. The loss function for DPO is structured as follows:
$\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}\right)\right]$ (10)
where $\sigma$ is the logistic function, $\beta$ is a parameter controlling the deviation from the base reference policy $\pi_{ref}$, namely the initial model $\pi^{LLM}$. In practice, the language model policy $\pi_{\theta}$ is also initialized to $\pi^{LLM}$. For one answer, denoted as $y_{w}\succ y_{l}|x$ where $y_{w}$ and $y_{l}$ denotes the preferred and dispreferred completion amongst $(y1,y2)$ respectively.

### v2

@@ 3.3 Ablation Study
As shown in Table 4, we conduct ablation experiments to evaluate the key components of LLaMA-Berry, using the solved rate of problems (Luo et al., 2024a) as a metric across benchmarks of increasing difficulty: GSM8K and AIME2024. Zero-shot CoT represents the base model’s greedy mathematical reasoning capabilities. SR-MCTS without PPRM refers to a basic version of our method that uses self-evaluation as the reward instead of PPRM and EBC.
When comparing iterative Self-Refine with SR-MCTS, especially on the GSM8K dataset, the introduction of MCTS effectively mitigates the issue of solution degradation into suboptimal results caused by flawed critiques in iterative methods. With rollouts of 8 and 16, SR-MCTS improves the solved rate of problems by 3.1% and 3.2%, respectively. Furthermore, in comparing SR-MCTS without PPRM and SR-MCTS with PPRM, PPRM further boosts the solved rate of problems on the relatively easier GSM8K dataset. With rollouts of 8 and 16, the solved rate of problems is elevated by 13% and 14.7%, respectively.
Table 4: Ablation study for LLaMA-Berry framework. The PPRM and MCTS components are removed to evaluate their effectiveness. We use the solved rate of problems as the metric and the base model is LLaMA-3.1-8B-Instruct.
GSM8K AIME24
Zero-shot CoT 76.6 6.7
Rollouts 8 16 8 16
Iterative Self-Refine 78.9 78.2 6.7 6.7
SR-MCTS without PPRM 82.0 81.4 6.7 6.7
SR-MCTS + PPRM (Ours) 95.0 96.1 16.7 26.7
Notably, on the more challenging AIME2024 dataset, both iterative Self-Refine and SR-MCTS without PPRM demonstrate limited search efficiency. However, SR-MCTS with PPRM can significantly improve the solved rate of problems from 6.7% (2/30) to 16.7% (5/30) and 26.7% (8/30) with rollouts of 8 and 16, respectively. The results underscore the efficacy of combining the Self-Refine method with PPRM when addressing complex problems. The contrast between self-reward and PPRM underscores the importance of designing reward mechanisms that can more effectively guide the search process. PPRM provides a more holistic incentive to the model, thus fostering more effective problem-solving strategies.
@@ 3.4 Scaling Study
Figure 3: Scaling of inference-time rollouts
To explore the potentials and trends of the scaling with rollouts in inference-time, we depict the solved rate of problems with rollouts in three benchmarks with different difficulty levels. Analyzing the performance alongside Figure 3, the increment in the number of rollouts consistently enhances model performance across various benchmarks, and the extent of these improvements differs depending on the benchmark’s complexity and the base model’s reasoning capability. These curves underscore that the performance of the LLaMA-Berry framework benefits from scaling up rollouts during inference, similar to the observations in OpenAI (2024). However, there are ceiling limitations, as seen in the GSM8K dataset, which suggest that the base model’s capabilities in both reasoning and refinement play a crucial role in determining the overall performance.
4 Related Works
Reward models for reasoning. Reliable reward models (Kang et al., 2024; Wang et al., 2023; Havrilla et al., 2024; Lightman et al., 2023; Ma et al., 2023) can effectively distinguish desirable responses from undesirable ones, which is especially important in complex reasoning. The outcome reward model (ORM) are trained with the final results of the reasoning paths. As the rewards are determined by the final answers, ORM may suffer from coarse supervision and misalignment issues. In contrast, process reward model (PRM) provides step-wise reward signals that are easier to interpret and guide the models to follow the CoTs. Therefore, PRM is generally considered to be more effective (Lightman et al., 2023). However, the success of PRM relies on extensive manually annotated data (Luo et al., 2024a; Havrilla et al., 2024), which is time-consuming and still faces the challenge of the volatility of absolute reward scores. Pairwise Preference Reward Model (PPRM) in  LLaMA-Berry converts absolute scoring into preference prediction task, which then brings robust reward signals via EBC method.
@@ 4 Related Works
Tree search reasoning. Sampling diverse reasoning paths (Brown et al., 2024) has demonstrated its effectiveness in enhancing the probability of finding the correct answers. Self-Consistency (Wang et al., 2022) samples a complete path each time while tree search methods like Tree-of-Thought (ToT) (Yao et al., 2023) and Monte Carlo Tree Search (MCTS) (Chen et al., 2024a, b; Luo et al., 2024b; Feng et al., 2023; Xie et al., 2024; Xu, 2023; Liu et al., 2023; Tian et al., 2024; Ding et al., 2023) extend multiple steps to optimize step answers and ultimately obtain the optimal solution. Additionally, Self-Refine (Madaan et al., 2023a) method has become a recent focus. Self-verification (Gero et al., 2023; Weng et al., 2022) and rStar (Qi et al., 2024) utilize the inherent capabilities of the model to iteratively explore and refine answers. However, the performance of Self-Refine is typically constrained by the inherent capabilities of the model, especially for small language models (SLMs) with significantly weaker Self-Refine abilities (Madaan et al., 2023b). Zhang et al. (2024b) suggests that the mathematical reasoning abilities of LLMs can be enhanced by treating the refinement process as a directed acyclic graph (DAG) through multi-agent collaboration. In our approach, we combine MCTS with Self-Refine to explore potential solutions and a global win-loss matrix is then constructed in the form of a directed graph to calculate the final quantile scores.
@@ 5 Conclusion
This research addresses challenges in complex mathematical reasoning, particularly at the Olympiad level, by enhancing the accuracy and efficiency of the search for reasoning paths. By introducing Self-Refine applied to Monte Carlo Tree Search (SR-MCTS), the LLaMA-Berry framework significantly improves the efficiency of solution generation by LLMs. Additionally, the Pairwise Preference Reward Model (PPRM) constructs preferences between solutions rather than merely scoring outcomes, calculating the final global quantile score using the enhanced Borda Count (EBC) method. Evaluation results demonstrate that LLaMA-Berry outperforms baseline approaches on benchmarks like GSM8K and MATH, and achieves competitive performance compared to closed-source models on Olympiad-level benchmarks such as AIME2024.
Limitation
The LLaMA-Berry framework has demonstrated strong performance in reasoning tasks, but there are still some challenges in practical applications. First, methods such as Monte Carlo Tree Search (MCTS) and Self-Refine have high computational costs. These techniques demand significant computational resources, which may limit the feasibility of deployment in environments with constrained computational capacity. As for summarizer of solutions, rule-based heuristics methods like self-consistency, major voting and mutual reasoning have shown a constraint to the ceiling search performance of MCTS. Thus, we aim to develop a learning-based summarizer as  OpenAI (2024) does to further enhance the search efficiency.
Furthermore, the current evaluation of the LLaMA-Berry framework has primarily focused on mathematical reasoning benchmarks, resulting in a relatively narrow assessment scope. As a result, its performance in broader domains, such as general knowledge, symbolic logic tasks, and multimodal applications, has not been sufficiently validated. In future work, we aim to improve the framework by evaluating it on a more diverse set of tasks to enhance its applicability.
Lastly, most experiments so far have utilized relatively small open-source models, with limited testing on larger or closed-source models. In future research, we plan to investigate the performance of LLaMA-Berry on larger models, particularly addressing challenges related to scaling and performance optimization.
@@ Appendix A Hyper-paramerer settings
All hyper-parameter settings for Section 3 are listed in Table A1.
Table A1: Hyper-parameter settings
@@ Appendix B Details of Grading and Metrics
Grading. We follow the correctness evaluation method proposed in PRM800K (Lightman et al., 2023) to score the answers generated by the algorithm. For the mathematical solutions proposed by the algorithm and their corresponding ground truth, we inform the model of the expected response format in a prompt. The answer’s formula or value is extracted by matching the response format with predefined rules. If the model fails to follow the expected format in the prompt and the rule-based extraction fails, the solution is directly judged as inconsistent with ground truth.
@@
    Following the criterion proposed in PRM800K (Lightman et al., 2023), we use the Sympy library (Meurer et al., 2017) to simplify the difference between the expression corresponding to the answer label, denoted as $a$, and the expression corresponding to the ground truth label, denoted as $b$. If the simplification yields $a-b\iff 0$, the criterion is satisfied.
Metrics. To provide a comprehensive and robust evaluation metric, we adopt major@k and rm@k as the evaluation metrics.
  • rm@k represents reward model best-of-N among $k$ sampled response (Yang et al., 2024c).
  • major@k (Kuncheva, 2014), also abbreviated as maj@k, is defined as the fraction of tasks where a majority of the top $k$ samples generated return the same correct solution. This metric focuses on consistency across multiple generated answers. The majority weight is calculated on the solution’s extracted labels.
  • The solved rate of problems (Lightman et al., 2023; Luo et al., 2024a) refers to the percentage of problems who have solutions meet the evaluation criteria.
For results without a subscript, the score represents the greedy evaluation using the default prompt of the base model. Results with a notation indicate the use of the corresponding prompt engineering technique, and those marked with self-consistency use the self-consistency aggregation method.
@@ C.1 Overview
The design goal of the PPRM is to integrate the properties of both PRM and ORM, providing a more nuanced preference prediction between any two solution answers. We attempt to utilize reinforcement learning methods for training, leveraging the model’s instruction-following capability to predict the relative merits of pairs of problem-solving answers. This will further enable the use of the EBC method to evaluate the global quantile scores of mathematical solutions.
@@ C.2 Data Collection
Figure A1: Dataset Construction of PPRM.
Our data synthesis is derived from two datasets: PRM800K (Lightman et al., 2023) and OpenMathInstruct-1 (Toshniwal et al., 2024). The PRM800K dataset, collected from the MATH dataset, comprises a substantial number of step-divided problem-solving answers, with manual quality annotations for each step. We primarily utilize this dataset to generate pairs of answers for comparative analysis based on step-wise process quality. The OpenMathInstruct-1 dataset incorporates data from the GSM8K and MATH datasets, which have been manually annotated for outcome correctness. We use this dataset to synthesize pairs for comparative analysis based on outcome quality.
In processing PRM800K, for a given problem, we first sample steps of varying quality annotations from the step-wise dataset to construct a complete reasoning path. For pairs with the same final step annotation, paths composed of higher-quality steps are considered superior to those with lower-quality steps. In cases where the final step annotations are not identical, the reasoning path with the superior final step annotation is regarded as better.
@@
Notably, we filtered out any data samples from PRM800K and OpenMathInstruct-1 that may overlap with the GSM8K and MATH test sets, especially the PRM800K. Ultimately, we formed a dataset of 7,780,951 entries for training the PPRM model.
@@ C.3 Direct Preference Pair Construction
For all pairs, we frame the inquiry as "For Question $Q$, is solution $a_{1}$ better than solution $a_{2}$?" If solution $a_{1}$ is deemed superior to solution $a_{2}$, we label it as ’Yes’; otherwise, it is labeled as ’No’.
@@ C.4 RLHF Training
Figure A2: RLHF training of PPRM.
We apply the DPO method to train the Gemma2-2B-it model using RLHF which is shown in Figure A2. The loss function for DPO is structured as $\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})=-\mathbb{E}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}\right)\right]$, where $\sigma$ is the logistic function, $\beta$ is a parameter controlling the deviation from the base reference policy $\pi_{ref}$, namely the initial model $\pi^{LLM}$. In practice, the language model policy $\pi_{\theta}$ is also initialized to $\pi^{LLM}$. For one answer, denoted as $y_{w}\succ y_{l}|x$ where $y_{w}$ and $y_{l}$ denote the preferred and dispreferred completion amongst $(\hat{y},y)$ respectively.